A square vegetable garden is to be tilled and then enclosed with a fence. If the fence costs $1 per foot and the cost of preparing the soil is $0.50 per ft^2, determine the size of the garden that can be enclosed for $120.
Thanks

Answer by scott8148(6628)

let x="length of a side" __ (.5)x^2+(1)4x=120

subtracting 120 and multiplying by 2 __ x^2+8x-240=0

factoring __ (x+20)(x-12)=0

x+20=0 __ x=-20 __ negative value not realistic

x-12=0 __ x=12